OpenStudy (anonymous):

The two spheres above have the same center. One has a radius of 4 cm, and the other has a radius of 5 cm. Approximately how much space is in between the two spheres?

OpenStudy (michele_laino):

the requested space is given by subtracting the volume of the littlest sphere, from the bigger one

OpenStudy (michele_laino):

if the radius of the littest sphere is r=4 cm, and the radius of the bigger sphere is R=5 cm, then the requested volume V is: \[\Large V = \frac{{4\pi }}{3}{R^3} - \frac{{4\pi }}{3}{r^3} = \frac{{4\pi }}{3}{5^3} - \frac{{4\pi }}{3}{4^3} = ...c{m^3}\]

OpenStudy (mrnood):

it asks 'approximately' so you could approximate pi to 3 so volume ~= 4(5^3-4^3)

OpenStudy (anonymous):

20

OpenStudy (mrnood):

@kaleamarie18 a) that is not correct b) do not give direct answers to questions (even if they are correct) You should guide the asker to the answer there is enough information above for him to work out the answer for himself

OpenStudy (mrnood):

@kaleamarie18 and BTW - welcome to Open Study